SKETCH THE GRAPH of f(x)= (3(x^2-4))/x^2-1 give the X and Y INTERCEPTS and the HORIZONTAL and VERTICAL ASYMPTOTES.
Answer by stanbon(75887)

SKETCH THE GRAPH of f(x)= (3(x^2-4))/x^2-1 give the X and Y INTERCEPTS and the HORIZONTAL and VERTICAL ASYMPTOTES.
------
x-int: y is zero when 3(x^2-4) = 0; x = 2 or x = -2
----
y-int: Let x = 0, then y = (3*-4)/(-1) = 12
----
Horizontal Asymptote(s): y = (3x^2)/x^2 = 3
---
Vertical Asymptote(s) when x^2-1 = 0; x = 1 or x = -1
============
Cheers,
Stan H.
===============